# Learning Geospatial Region Embedding with Heterogeneous Graph

### v1

@@ Appendix A More Introduction of ESA WorldCover Dataset
WorldCover dataset is continuously updated and revised by ESA with a highly efficient processing pipeline and scalable infrastructure. The core model is trained on a total of 2,160,210 Sentinel-2 images and is able to process the whole world in less than 5 days [49, 50]. The revised version we utilized in this paper was released on 28 October 2022, which elevated the global overall accuracy to 76.7% [50]. This version is free of charge to the entire community and widely accepted by the United Nations Convention to Combat Desertification (UNCCD), the World Resources Institute (WRI), the Centre for International Forestry Research (CIFOR), the Food and Agriculture Organization (FAO) and the Organisation for Economic Co-operation and Development (OECD). We visualized 3 examples of the WorldCover data unitized in this paper in Figure 9.
@@
Appendix B Motivation for Using Entity Segmentation for Satellite Imagery Encoding
B.1 Introduction of semantic segmentation-based approach
It is noticeable that conventional vision encoders are devised and trained on natural images, which significantly differ from satellite imagery. It is a formidable task to comprehend the intricate and specialized geo-semantic content present within satellite imagery.
Fortunately, satellite images exhibit a high degree of structural organization in terms of semantic information, unlike natural images that contain a myriad of diverse information. To effectively interpret a satellite image, one only needs to concentrate on the geological entities it encompasses, the spatial extent of these entities, and their respective positions within the geospace. Consequently, it is highly promising to employ an entity segmentation approach for encoding satellite imagery, as it can directly provide us with essential information regarding the entities and their spatial coverage.
In our proposed method, GeoHG, we employ an entity segmentation-based framework as the backbone for encoding satellite imagery, as opposed to the presently prevalent vision encoders, such as CNN and Transformer. To implement this approach, we utilize the ESA WorldCover Dataset, as discussed in Appendix A, due to its extensive validation by geographical researchers and demonstrated robust global generalization capabilities. As depicted in Figure 10, we directly retrieve the geo-entity segmentation results from the WorldCover Dataset for the input satellite imagery. Subsequently, we determine the proportion of geo-entities within the relevant region and construct the environmental feature embedding $E_{env}$.
Figure 10: Geo-entity Segmentation for Satellite Images based on WorldCover Dataset.
@@ B.1 Introduction of semantic segmentation-based approach
B.2 Validation experiment for segmentation-based approach
To validate the effectiveness of our segmentation-based approach. We devise a validation experiment on predicting five indicators for 4 cities by taking the segmentation-based geo-entity proportion $E_{env}$ as the only input (named GeoSegment for convenience in Tabel 2). In comparison to the state-of-the-art LLM-enhanced satellite imagery encoder UrbanCLIP [42], POI information-enhanced imagery encoder PG-SimCLR [38] and two traditional image encoders ResNet-18 and AutoEncoder. These baselines are detailed and introduced in Appendix E.
@@ B.2 Validation experiment for segmentation-based approach
[…]
From Table 2, it is evident that our semantic segmentation-based approach overwhelmingly outperforms all baseline models in almost every task. This includes the state-of-the-art LLM-enhanced satellite imagery encoder, UrbanCLIP [42], and the POI information-enhanced encoder, PG-SimCLR [38]. Notably, GeoSegment achieves these significant advantages without relying on any additional information, such as LLM descriptions or POI data. This demonstrates the superiority of our method in satellite imagery encoding.
Appendix C More Details of GeoHG
C.1 Basic graph theory and our motivation
Graph theory. Graphs provide a natural abstraction to represent structured relationships between entities (nodes) and their attributes (features). Within this framework, knowledge and information are organized via connectivity patterns encoded by edges. A key advantage of graph-structured representations is that nodes can reason about their representations not just from their attributes, but also by recursively aggregating information from their neighbors. This allows graphs to effectively capture complex relational dependencies and contextual patterns.
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for learning representations on graph-structured data by leveraging the message-passing mechanism. The core idea of this mechanism involves each node recursively aggregating representation vectors from its local neighborhood, allowing it to accumulate information from an expanding neighborhood scope across iterations [37]. Formally, we can represent this process through the message-passing equation:
@@ C.1 Basic graph theory and our motivation
Hypergraph theory. Traditional graph models can only connect pairs of vertices within edges, thereby facing limitations in representing higher-order multi-way relationships beyond simple pairwise associations. In real-world scenarios, relationships between data typically go beyond simple pairwise links and involve complex multi-element patterns. For example, in a transportation network, a route often spans multiple cities or locations rather than just directly connecting two endpoints, which is difficult to capture using traditional graphs. To overcome this, hypergraph theory generalizes the graph formulation by introducing hyperedges that can connect any number of vertices, naturally facilitating the representation of higher-order multi-way relationships, defined as:
$G=(V,E)\quad where\quad V=\{v_{1},v_{2},...,v_{N}\},E=\{\mathcal{H}_{1},\mathcal{H}_{2},...,\mathcal{H}_{M}\}$ (12)
@@
Motivations for modeling geospatial representations with hypergraph. In the context of geospatial representation modeling, the relationships between geospatial areas go far beyond simple spatial proximity, instead arising from the intricate interplay of various environmental and societal factors. Climatic conditions, topography, population distribution, economic factors, and more, exhibit intricate high-order influence patterns across different regions. Traditional methods fail to explicitly characterize this inherent high-order relational structure [7, 59]. Although multi-view graphs for geospace are capable of incorporating additional connections between regions by introducing extra edges between graph nodes, this approach comes with considerable complexity due to its inefficient representation. Furthermore, it is questionable whether distant regions are directly influencing each other just like their connections in a multi-view graph. For instance, normally, a region might initially deteriorate the overall water body of the geospace, subsequently affecting a remote region with abundant water resources, rather than through direct connections.
To effectively model such higher-order geospatial relations, we treat each geospatial region as a node and abstract the higher-order environmental/societal associations influencing multiple regions as hyperedges connecting all the relevant nodes, precisely capturing these intricate dependency patterns. Building upon this idea, we further propose GeoHG to encode high-order relationships and intricate characteristics between regions.
C.2 Details of GeoHG structure
GeoHG introduces a novel heterogeneous hypergraph representation, termed GeoHG, to encode the high-order relational structures within geospatial data. GeoHG comprises three types of nodes: regional nodes, environmental entity nodes, and societal entity nodes, along with the relations between them. Specifically, regional nodes correspond to $1km\times 1km$ grid cells from the geographic space; environmental entity nodes represent the 9 distinct geo-entity classes detected from satellite imagery; societal entity nodes constitute the 14 different point-of-interest (POI) categories within the regions.
We construct hyperedges connecting multiple regional nodes through the environmental and societal entity nodes, enabling the explicit modeling of higher-order inter-region relationships induced by environmental or societal factors. The semantics of the node types and their relationships within GeoHG are illustrated in Tables 3 and 4, respectively.
Table 3: Major entities in GeoGraph
@@ C.2 Details of GeoHG structure
Table 4: Major relations in GeoGraph
@@
Appendix D Dataset and Experiment Settings
Dataset Details. We employ five representative tasks located in four cities in China: Beijing, Shanghai, Guangzhou and Shenzhen. Population, GDP, and Night Light tasks reflect anthropogenic activities, whereas Carbon and Temperature tasks characterize the natural environment. We conduct data preprocessing according to [42]. The detailed information of datasets is listed below:
@@ Appendix D Dataset and Experiment Settings
  • $PM_{2.5}$: The $PM_{2.5}$ dataset is sourced from ChinaHighPM2.5 dataset [36]. This dataset combines ground-based observations, atmospheric reanalysis, emission inventory, and other techniques to obtain nationwide seamless ground PM2.5 data from 2000 to the present. The main scope is the entire China area, the spatial resolution is 1 km, the time resolution is daily, monthly, and yearly, and the unit is µg/m3.
Table 5: Dataset statistics.
Dataset Coverage in Geospace Satellite Image POI Information
Bottom-left Top-right Area ($km^{2}$)
Beijing 39.75°N, 116.03°E 40.15°N, 116.79°E 4,277 4,277 709,232
Shanghai 30.98°N, 121.10°E 31.51°N, 121.80°E 5,292 5,292 808,957
Guangzhou 22.94°N, 113.10°E 23.40°N, 113.68°E 8,540 8,540 805,997
Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 5,150 5,150 717,461
Shenzhen-enlarged 22.45°N, 113.75°E 23.84°N, 114.62°E 17,264 17,264 1,813,547
Implementation Details. We implement GeoHG, GeoHG-SSL and baselines with PyTorch 3.9 on a single NVIDIA RTX A6000 with 24GB relevant memory. We use a 2-layer MLP as the regression head for prediction. GeoHG is trained using Adam optimizer with a learning rate of 0.01. For the hyperedge gate $\theta_{Env}$ and $\theta_{Soc}$, we conduct grid searches over (0.2, 0.4, 0.6, 0.8) and (0.3,0.6,0.9,1.2,1.5) respectively. For the number of layers of the graph convolutional block, we test it from 1 to 3. The execution time required for training the GeoHG is approximately 3 minutes per task.
Final Settings of GeoHG. We introduce the best hyperparameter configurations for each task as below:
@@
Evaluation Metrics. We employ Mean Absolute Error (MAE), rooted mean squared error (RMSE), and coefficient of determination ($R^{2}$) as evaluation metrics, these metrics are calculated as below:
@@
Appendix E Details of Baselines
Description & Settings. We have chosen to incorporate a variety of widely utilized traditional methods and prominent cutting-edge techniques for comparative evaluation. We select two classical methods (AutoEncoder [18] and ResNet-18 [10]), four state-of-the-art vision-based models (Urban2Vec [35] and UrbanCLIP [42]) and multi-modal approaches (UrbanVLP [9] and GeoStructual [20]) for geospatial region embedding. Moreover, we explore a variation of GeoHG where we employ Graph Self-Supervised Learning (SSL) pretraining instead of end-to-end training, represented as GeoHG-SSL. We describe these baselines as follows.
  • AutoEncoder [18]: A neural network architecture framework targets unlabeled satellite imagery for feature extraction using reconstruction loss.
  • ResNet-18 [10] The renowned residual neural network, pretrained on the extensive ImageNet dataset [2], is capable of directly extracting visual features from satellite imagery by leveraging the knowledge it has previously acquired from natural imagery.
@@ Appendix E Details of Baselines
  • UrbanCLIP: [42] A model incorporating a multi-modal Large Language Model (LLM) to enhance the encoding of satellite imagery. This model generates descriptive texts for satellite images using the LLM and then fuses these texts with the images through an image-text contrastive learning-based approach to capture the complexity and diversity of geospatial areas.
@@
  • PG-SimCLR [38]: A contrastive learning framework that introduces societal information (i.e., POI) into geospatial region representation learning from satellite imagery.
@@
  • GeoStructural [20]: The graph-based framework profiles geospatial regions by utilizing street segments as graph structure for adaptively fusing features from multi-level satellite and street-view images. For convenience, we refer to this method as GeoStructural.
@@
  • UrbanVLP [9]: An region embedding method based on contrastive learning, which incorporates satellite imagery, street-view images, and spatial position structure. This method is further enhanced by incorporating a Large Language Model (LLM) and GeoCLIP [33], resulting in improved robustness and performance.
Appendix F More Details about Experiments Results
F.1 RMSE metrics of experiment results
Given the page constraints, we hereby present the complementary version of our experimental results table in Table 6, which includes RMSE results for each experiment. It is discernible that the RMSE outcomes are consistent with other performance metrics, thereby conclusively affirming the overall superiority of our model’s performance.
Table 6: Region indicators prediction results. The bold/underlined font means the best/the second-best result.
Methods GeoHG GeoHG-SSL UrbanVLP GeoStructural PG-SimCLR UrbanCLIP ResNet-18 AutoEncoder
Metric $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE $\text{R}^{2}$ RMSE
Beijing Carbon 0.954 0.201 0.937 0.224 0.787 0.457 0.765 0.472 0.442 0.754 0.664 0.598 0.394 0.805 0.298 0.844
Population 0.874 0.351 0.870 0.374 0.725 0.513 0.730 0.504 0.471 1.117 0.461 0.669 0.266 0.857 0.168 0.918
GDP 0.647 0.567 0.644 0.581 0.586 0.650 0.617 0.609 0.277 1.254 0.355 0.864 0.285 1.024 0.171 1.251
Night Light 0.901 0.311 0.900 0.353 0.531 0.629 0.488 0.657 0.369 0.728 0.420 0.700 0.348 0.734 0.276 0.784
$PM_{2.5}$ 0.971 0.160 0.970 0.169 0.641 0.594 0.694 0.482 0.398 0.845 0.533 0.699 0.341 0.821 0.209 0.840
Shanghai Carbon 0.915 0.290 0.912 0.312 0.716 0.529 0.688 0.557 0.298 0.914 0.671 0.569 0.326 0.763 0.230 0.771
Population 0.936 0.244 0.928 0.251 0.593 0.607 0.613 0.583 0.315 0.959 0.456 0.748 0.279 0.797 0.166 0.898
GDP 0.778 0.468 0.767 0.477 0.310 0.816 0.377 0.721 0.294 1.052 0.326 0.807 0.289 1.137 0.197 1.511
Night Light 0.898 0.311 0.891 0.347 0.457 0.667 0.442 0.685 0.308 0.768 0.387 0.709 0.244 0.756 0.164 0.729
$PM_{2.5}$ 0.866 0.379 0.836 0.394 0.486 0.654 0.527 0.592 0.303 0.895 0.444 0.774 0.292 0.954 0.243 0.942
Guangzhou Carbon 0.885 0.336 0.884 0.371 0.698 0.514 0.681 0.529 0.422 0.708 0.585 0.603 0.375 0.673 0.254 0.733
Population 0.871 0.244 0.855 0.255 0.665 0.441 0.687 0.433 0.303 0.972 0.533 0.687 0.274 0.831 0.195 0.928
GDP 0.715 0.532 0.712 0.574 0.436 0.764 0.439 0.699 0.282 1.264 0.440 0.762 0.251 1.048 0.176 1.463
Night Light 0.871 0.378 0.854 0.391 0.577 0.573 0.574 0.581 0.435 0.627 0.483 0.633 0.242 0.704 0.176 0.798
$PM_{2.5}$ 0.833 0.403 0.822 0.414 0.638 0.624 0.652 0.597 0.315 0.741 0.56 0.694 0.231 0.859 0.196 0.903
Shenzhen Carbon 0.926 0.290 0.912 0.304 0.659 0.568 0.647 0.581 0.257 0.816 0.562 0.571 0.241 0.726 0.189 0.887
Population 0.892 0.244 0.879 0.261 0.790 0.448 0.797 0.390 0.311 0.892 0.527 0.610 0.299 0.855 0.175 0.921
GDP 0.798 0.468 0.767 0.489 0.532 0.676 0.517 0.699 0.307 1.003 0.508 0.693 0.234 1.125 0.119 1.733
Night Light 0.942 0.245 0.939 0.268 0.457 0.667 0.445 0.682 0.454 0.588 0.387 0.709 0.243 0.719 0.166 0.758
$PM_{2.5}$ 0.906 0.298 0.905 0.331 0.566 0.598 0.597 0.491 0.323 0.883 0.430 0.762 0.273 0.826 0.149 0.891
F.2 Mean and standard deviation of metrics for GeoHG
Each method is executed five times, and we report the detailed mean and standard deviation of both metrics for GeoHG in Table 7.
Table 7: The mean and standard deviation of both metrics in 5-run for GeoHG.
Cities Beijing Shanghai
Metric $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE
Carbon 0.954±0.002 0.11±0.002 0.201±0.002 0.915±0.001 0.157±0.001 0.29±0.007
Population 0.874±0.003 0.271±0.001 0.351±0.001 0.936±0.005 0.161±0.001 0.244±0.001
GDP 0.647±0.015 0.336±0.010 0.567±0.009 0.778±0.001 0.323±0.002 0.468±0.001
Night Light 0.901±0.014 0.239±0.001 0.311±0.002 0.898±0.004 0.222±0.003 0.311±0.004
$PM_{2.5}$ 0.971±0.009 0.064±0.002 0.16±0.003 0.866±0.001 0.12±0.002 0.3789±0.003
Cities Guangzhou Shenzhen
Metric $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE
Carbon 0.885±0.001 0.219±0.001 0.336±0.001 0.926±0.001 0.128±0.001 0.29±0.001
Population 0.871±0.003 0.244±0.005 0.368±0.002 0.892±0.005 0.165±0.007 0.244±0.005
GDP 0.715±0.002 0.371±0.005 0.532±0.002 0.798±0.002 0.297±0.002 0.468±0.003
Night Light 0.871±0.001 0.234±0.002 0.378±0.001 0.942±0.001 0.149±0.001 0.245±0.001
$PM_{2.5}$ 0.833±0.003 0.133±0.002 0.4034±0.003 0.906±0.002 0.116±0.002 0.298±0.004
F.3 More ablation study results
The ablation results for Beijing and Shenzhen are illustrated in Figure 11. GeoHG-Mono only preserves the isomorphic graph structure and adjacency relationships between regions, while GeoHG-MLP completely discards the graph structure. Similar to the results for Guangzhou and Shanghai in Figure 5, discarding the heterogeneous graph structure leads to severe performance degradation across all tasks. Furthermore, removing environmental, social, or location information also reduces model performance, but the degree of degradation varies across different cities when omitting each type of information.
Figure 11: Results of Ablation Study on $R^{2}$ Metric for Beijing and Shenzhen.
F.4 Details of data efficiency experiment
Table 8: Region indicators prediction results in the few-shot setting. The evaluation metric is $R^{2}$, and the test set constitutes a 20% random sample disjoint from the training data. The bold/underlined font means the best/the second-best result.
Methods GeoHG UrbanVLP GeoStructural UrbanCLIP
Available Data 20% 10% 5% 20% 10% 5% 20% 10% 5% 20% 10% 5%
[…]
We provide the performances of several baseline models in a data-limited setting, as shown in Table 8. We observed that existing geospatial embedding models exhibit certain few-shot learning capabilities, performing relatively well despite significantly reduced training data. Our model consistently outperforms other models under the same training data configurations. Moreover, our model can achieve performance exceeding or comparable to that of other models while using less data.
F.5 Details of investigation of GeoHG with GNNExplainer
To validate GeoHG’s efficacy in capturing higher-order geospatial relationships, we employ GNNExplainer [46], a model-agnostic technique for interpreting GNN predictions through identifying crucial subgraph structures and features. Specifically, for the carbon emission prediction task on the region #(35,65) in Guangzhou, we utilized GNNExplainer to extract the top 10 most influential edges, as shown in Table 9. Similar results for region #(43,87) are shown in Table 10. Based on Table  9, we can derive the following key findings:

### v2

@@ Appendix B More Introduction of ESA WorldCover Dataset
The ESA WorldCover dataset  [^1] https://esa-worldcover.org/en represents a groundbreaking advancement in land use/land cover mapping, offering freely accessible, high-resolution (10 m) global coverage based on satellite imagery. Inspired by the 2017 WorldCover conference  [^2] https://worldcover2017.esa.int/, the European Space Agency (ESA) launched the WorldCover project and the primary accomplishment of this endeavor was the introduction in October 2021 of a freely accessible global WorldCover Dataset at a groundbreaking 10 m resolution for the year 2020 [Zanaga et al., 2021, 2022]. This dataset leverages satellite imagery from both Sentinel-1 and Sentinel-2, and encompasses 11 distinct geo-entity categories, shown in Figure 9. It has also undergone rigorous independent validation by Wageningen University (for statistical accuracy) and the International Institute for Applied Systems Analysis (IIASA) (for spatial accuracy), attaining a notable global overall accuracy of approximately 75% [Venter et al., 2022].
@@
Figure 11: Geo-entity Segmentation for Satellite Images based on ESA WorldCover.
WorldCover dataset is continuously updated and revised by ESA with a highly efficient processing pipeline and scalable infrastructure. The core model is trained on a total of 2,160,210 Sentinel-2 images and is able to process the whole world in less than 5 days [Zanaga et al., 2021, 2022]. The revised version we utilized in this paper was released on 28 October 2022, which elevated the global overall accuracy to 76.7% [Zanaga et al., 2022]. This version is free of charge to the entire community and widely accepted by the United Nations Convention to Combat Desertification (UNCCD), the World Resources Institute (WRI), the Centre for International Forestry Research (CIFOR), the Food and Agriculture Organization (FAO) and the Organisation for Economic Co-operation and Development (OECD). We visualized 3 examples of the WorldCover data unitized in this paper in Figure 10.
Appendix C Entity Segmentation for Satellite Image Encoding
C.1 Semantic segmentation-based approach
It is apparent that traditional vision encoders are designed and trained on natural images, which are considerably different from satellite images. Understanding the complex and highly specialized geo-semantic information in satellite imagery is a challenging task.
Luckily, satellite images display a significant level of structured organization concerning semantic content, in contrast to natural images, which contain a vast array of diverse information. To accurately analyze a satellite image, one needs to focus solely on the geological entities it includes, the extent of these entities, and their positions within the geospace. Therefore, using an entity segmentation method to encode satellite images is very promising, as it can directly offer crucial information regarding the entities and their spatial distribution.
In our proposed method, GeoHG, we employ an entity segmentation-based framework as the backbone for encoding satellite imagery, as opposed to the presently prevalent vision encoders, such as CNN and Transformer. To implement this approach, we utilize the ESA WorldCover Dataset, as discussed in Appendix B, due to its extensive validation by geographical researchers and demonstrated robust global generalization capabilities. As depicted in Figure 11, we directly retrieve the geo-entity segmentation results from the WorldCover Dataset for the input satellite imagery. Subsequently, we determine the proportion of geo-entities within the relevant region and construct the environmental feature embedding $E_{env}$.
@@ C.1 Semantic segmentation-based approach
PG-SimCLR 0.257 0.683 0.816 0.311 0.758 0.892 0.307 0.895 1.003 0.454 0.488 0.682 0.597 0.451 0.491
UrbanCLIP 0.562 0.483 0.571 0.527 0.592 0.610 0.508 0.464 0.693 0.387 0.511 0.709 0.460 0.586 0.762
GeoSegment 0.542 0.518 0.698 0.602 0.438 0.602 0.658 0.411 0.587 0.738 0.386 0.512 0.471 0.480 0.752
C.2 Validation experiment
To validate the effectiveness of our segmentation-based approach. We devise a validation experiment on predicting five indicators for 4 cities by taking the segmentation-based geo-entity proportion $E_{env}$ as the only input (named GeoSegment for convenience in Tabel 4). In comparison to the state-of-the-art LLM-enhanced satellite imagery encoder UrbanCLIP [Yan et al., 2024], POI information-enhanced imagery encoder PG-SimCLR [Xi et al., 2022] and two traditional image encoders ResNet-18 and AutoEncoder. These baselines are detailed and introduced in Appendix F.
From Table 4, it is evident that our semantic segmentation-based approach overwhelmingly outperforms all baseline models in almost every task. This includes the state-of-the-art LLM-enhanced satellite imagery encoder, UrbanCLIP [Yan et al., 2024], and the POI information-enhanced encoder, PG-SimCLR [Xi et al., 2022]. Notably, GeoSegment achieves these significant advantages without relying on any additional information, such as LLM descriptions or POI data. This demonstrates the superiority of our method in encoding satellite images.
Appendix D Theoretical Analysis and Details of GeoHG
D.1 Graph theory and our motivation
D.1.1 Basic graph theory.
Graphs provide a natural abstraction to represent structured relationships between entities (nodes) and their attributes (features). Within this framework, knowledge and information are organized via connectivity patterns encoded by edges. A key advantage of graph-structured representations is that nodes can reason about their representations not just from their attributes, but also by recursively aggregating information from their neighbors. This allows graphs to effectively capture complex relational dependencies and contextual patterns.
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for learning representations on graph-structured data by leveraging the message-passing mechanism. The core idea of this mechanism involves each node recursively aggregating representation vectors from its local neighborhood, allowing it to accumulate information from an expanding neighborhood scope across iterations [Wu et al., 2020]. Formally, we can represent this process through the message-passing equation:
@@ D.1 Graph theory and our motivation
D.1.2 Hypergraph theory.
Traditional graph models can only connect pairs of vertices within edges, thereby facing limitations in representing higher-order multi-way relationships beyond simple pairwise associations. In real-world scenarios, relationships between data typically go beyond simple pairwise links and involve complex multi-element patterns. For example, in a transportation network, a route often spans multiple cities or locations rather than just directly connecting two endpoints, which is difficult to capture using traditional graphs. To overcome this, hypergraph theory generalizes the graph formulation by introducing hyperedges that can connect any number of vertices, naturally facilitating the representation of higher-order multi-way relationships, defined as:
$G=(V,E);V=\{v_{1},v_{2},...,v_{N}\},E=\{\mathcal{H}_{1},\mathcal{H}_{2},...,\mathcal{H}_{M}\}$ (12)
@@
D.2 Details of GeoHG structure
Table 5: Major entities in GeoHG
@@ D.2 Details of GeoHG structure
Table 6: Major relations in GeoHG
@@
We introduce a novel heterogeneous hypergraph representation, termed GeoHG, to encode the high-order relational structures within geospatial data. GeoHG comprises three types of nodes: regional nodes, environmental entity nodes, and societal entity nodes, along with the relations between them. Specifically, regional nodes correspond to $1km\times 1km$ grid cells from the geographic space; environmental entity nodes represent the 9 distinct geo-entity classes detected from satellite imagery; societal entity nodes constitute the 14 different point-of-interest (POI) categories within the regions.
We construct hyperedges connecting multiple regional nodes through the environmental and societal entity nodes, enabling the explicit modeling of higher-order inter-region relationships induced by environmental or societal factors. The semantics of the node types and their relationships within GeoHG are illustrated in Tables 5 and 6, respectively.
Appendix E Dataset and Experiment Settings
E.1 Dataset details.
We employ five representative tasks located in four cities in China: Beijing, Shanghai, Guangzhou and Shenzhen. Population, GDP, and Night Light tasks reflect anthropogenic activities, whereas Carbon and Temperature tasks characterize the natural environment. We conduct data preprocessing according to [Yan et al., 2024]. The detailed information of datasets is listed below:
@@ Appendix E Dataset and Experiment Settings
  • $PM_{2.5}$: The $PM_{2.5}$ dataset is sourced from ChinaHighPM2.5 dataset [WEI Jing, 2024]. This dataset combines ground-based observations, atmospheric reanalysis, emission inventory, and other techniques to obtain nationwide seamless ground PM2.5 data from 2000 to the present. The main scope is the entire China area, the spatial resolution is 1 km, the time resolution is daily, monthly, and yearly, and the unit is µg/m3.
E.2 External experiment settings of GeoHG.
We introduce the best hyperparameter configurations for each task as below:
@@
E.3 Evaluation metrics.
We employ Mean Absolute Error (MAE), rooted mean squared error (RMSE), and coefficient of determination ($R^{2}$) as evaluation metrics, these metrics are calculated as below:
@@
Appendix F Details of Baselines
F.1 Description & Settings
. We describe the baselines as follows:
  • IDW (Inverse Distance Weighting) [Zimmerman et al., 1999] A simple and widely used spatial interpolation method that estimates values at unsampled locations by weighting the values of nearby sampled points inversely by their distance to the target location. It assumes that the influence of a point decreases monotonously with distance, making it effective when data points are dense and evenly distributed.
  • Universal Kriging [Brus and Heuvelink, 2007] A geostatistical interpolation technique that considers both the distance and the spatial correlation between data points. It incorporates a trend component into the interpolation model, allowing for more accurate predictions in the presence of spatially varying trends.
@@ F.1 Description & Settings
  • UrbanCLIP: [Yan et al., 2024] A model incorporating a multi-modal Large Language Model (LLM) to enhance the encoding of satellite imagery. This model generates descriptive texts for satellite images using the LLM and then fuses these texts with the images through an image-text contrastive learning-based approach to capture the complexity and diversity of geospatial areas.
@@
  • PG-SimCLR [Xi et al., 2022]: A contrastive learning framework that introduces societal information (i.e., POI) into geospatial region representation learning from satellite imagery.
@@
  • GeoStructural [Li et al., 2022]: The graph-based framework profiles geospatial regions by utilizing street segments as graph structure for adaptively fusing features from multi-level satellite and street-view images. For convenience, we refer to this method as GeoStructural.
@@
  • UrbanVLP [Hao et al., 2024]: An region embedding method based on contrastive learning, which incorporates satellite imagery, street-view images, and spatial position structure. This method is further enhanced by incorporating a Large Language Model (LLM) and GeoCLIP [Vivanco Cepeda et al., 2024], resulting in improved robustness and performance.
F.2 GNNExplainer
To validate GeoHG’s efficacy in capturing higher-order geospatial relationships, we employ GNNExplainer [Ying et al., 2019], a model-agnostic technique for interpreting GNN predictions through identifying crucial subgraph structures and features. Specifically, for the carbon emission prediction task on the region #(35,65) in Guangzhou, we utilized GNNExplainer to extract the top 10 most influential edges, as shown in Table 7. Similar results for region #(43,87) are shown in Table 8. Based on Table  7, we can derive the following key findings: